OpenStudy (4meisu):

In an arithmetic sequence, the first term is 5 and the fourth term is 40. Find the second term.

OpenStudy (anonymous):

a1=5 a4=40 a4 = (4-1)d +a1 40 =3d +5 d= 35/3 a2 =a1 +d =5 + 35/3 = 50/3